OpenStudy (anonymous):

2x - y = 2s 2x - 2y = 4s The lines whose equations are shown intersect at which point?

OpenStudy (anonymous):

I believe you're going to solve this like any other system of equations. You can use the elimination method, eliminating the x. Since both equations have a 2x, multiply one of the two equations (it doesn't really matter which, but I would negate the second one) and then you can add the two equations together. The only difference is that you'll get an answer in terms of "s" instead of an actual number.

OpenStudy (anonymous):

im lost...

OpenStudy (anonymous):

Ok, I'm sure you've looked at solving systems of equations before. You can use the elimination method. Multiply the entire second equation by -1 and what do you get?

OpenStudy (anonymous):

-2x+2y=0?

OpenStudy (abb0t):

If you don't like those, you can use cramers rule, using matrices. a 3x2 matrix.

OpenStudy (anonymous):

The left side is right. But 4s times -1 is not 0. It's -4s. So now we have these equations, so we'll add them: 2x -y= 2s -2x +2y= -4s -------------- 0 + y = -2s y=-2s

OpenStudy (anonymous):

alright. Yea use actually never done this before. My teacher skipped over this unit

OpenStudy (anonymous):

Ok, well now you have y. So you can take either equation and plug in -2s for y. Then you can solve for x. Once you get x, you have the point.

OpenStudy (anonymous):

But wait would the coordinates be (-2s,0) or (0,-2s)

OpenStudy (anonymous):

Once you plugin you'll get x=0. Since x=0 and y=-2s, they intersect at the point (0, -2s) since the x coordinate goes first.

OpenStudy (anonymous):

alright thank you so much!

OpenStudy (anonymous):

You're welcome, glad I could help